# Determinantal random subgraphs

### v2

@@ 1.3.1. The connected case
where $\mathsf{T}(\underline{x})=\sum_{T\in\mathcal{T}(\mathsf{G})}\underline{x}^{T}$, often called the Kirchhoff polynomial, is the generating function of spanning trees of $\mathsf{G}$, the set of which we write $\mathcal{T}(\mathsf{G})$.
@@ 9.2. The case of two-dimensional simplicial complexes
For example, take $X$ to be a $2$-cellulation of a closed surface $\Sigma$ of genus $g\geqslant 1$. Then $b_{1}(X)=2g$ and $H_{1}(X,\mathbb{Z})\simeq H_{1}(\Sigma,\mathbb{Z})\simeq\mathbb{Z}^{2g}$. Elements of the support of our measure, $\mathcal{H}_{1}(X)$, are then sometimes called $g$-quasitrees of the map $X$ in the combinatorics literature. [^6] In yet other parts of the literature, $g$-quasitrees of a map $X$ of genus $g$ are also refered to as unicellular maps, that is maps whose $1$-skeleton is a subgraph of the $1$-skeleton of the map such that the complement of this subgraph in the map $X$ is homeomorphic to a disc. We can obtain determinantal measures on these lower genera quasi-trees by taking other subspaces $H$ such that $\mathop{\mathrm{im}}d\subseteq H\subseteq\ker d^{\prime}$. Along with lower genera quasitrees, they appear in the definition of the Bollobás–Riordan polynomial of the cellulation [CKS11], which is known to fit in the general framework of Tutte polynomials of matroids [MS18].
@@ 14.1.1. Determinantal cycle-rooted spanning forests
Forman [For93] proved a matrix-tree type formula for $\mathcal{U}(\mathsf{G})$ and this implies the existence of a determinantal probability measure on $\mathcal{U}(\mathsf{G})$ according to Proposition 5.3, a fact first proved by Kenyon [Ken11] who rediscovered Forman’s result, and extended it to the quaternion case. See Figure 9.
This determinantal probability measure is determined by an element $\theta\in\Omega^{1}(\mathsf{G})$, and the weight assigned to a cycle-rooted spanning forest is
(143) $w(F)\,\underline{x}^{F}=\underline{x}^{F}\prod_{c\,\text{cycle}}|1-\mathrm{hol}_{h}(c)|^{2}\,,$
where $\mathrm{hol}_{h}(c)=e^{i\theta(c)}$ is the holonomy of the cycle $c$. The corresponding subspace of $\Omega^{1}(\mathsf{G})$ is the range of a discrete covariant derivative $d_{h}$, a twisted analogue of the coboundary operator $d$, where $h$ is a connection which represents $\theta$ in the sense that $h_{e}=\exp(i\theta_{e})$ (for detailed definitions, see [Ken11, Kas15, KL22, KL25]).
@@
14.1.2. An example
@@
Let us consider a unitary rank $1$ connection $h$ on $\mathsf{G}$. As we have just seen, the subspace $\mathop{\mathrm{im}}d_{h}$ of $\Omega^{1}(\mathsf{G})$ defines a determinantal probability measure on $\mathcal{U}(\mathsf{G})$. Let this be our reference distribution.
@@
We consider now a $1$-form $\theta$ outside $\mathop{\mathrm{im}}d_{h}$ and consider the subspace $H_{\theta}=\mathop{\mathrm{im}}d_{h}\oplus\mathbb{C}\theta\subseteq\Omega^{1}(\mathsf{G})$. By Theorem 13.13, the corresponding determinantal measure is supported on the set of spanning subgraphs $K$ with connected components $K_{1},\ldots,K_{m}$, all of which have first Betti number equal to $1$, but $K_{1}$, which has first Betti number equal to $2$.

### v3

@@ 1.3.1. The connected case
where $\mathsf{T}(\underline{x})=\sum_{T\in\mathcal{T}(\mathsf{G})}\underline{x}^{T}$, often called the Kirchhoff polynomial, is the generating function of spanning trees of $\mathsf{G}$, the set of which we write $\mathcal{T}(\mathsf{G})$. In the case $k=1$, Equation (3) already appeared in [KW15, Lemma 2] and [KK17, Lemma 4].
@@ 9.2. The case of two-dimensional simplicial complexes
For example, take $X$ to be a $2$-cellulation of a closed surface $\Sigma$ of genus $g\geqslant 1$. Then $b_{1}(X)=2g$ and $H_{1}(X,\mathbb{Z})\simeq H_{1}(\Sigma,\mathbb{Z})\simeq\mathbb{Z}^{2g}$. Elements of the support of our measure, $\mathcal{H}_{1}(X)$, are then sometimes called $2g$-quasitrees of the map $X$ in the combinatorics literature. [^6] In yet other parts of the literature, $2g$-quasitrees of a map $X$ of genus $g$ are also refered to as unicellular maps, that is maps whose $1$-skeleton is a subgraph of the $1$-skeleton of the map such that the complement of this subgraph in the map $X$ is homeomorphic to a disc. We can obtain determinantal measures on these lower genera quasi-trees by taking other subspaces $H$ such that $\mathop{\mathrm{im}}d\subseteq H\subseteq\ker d^{\prime}$. An special example is given in Section 12.1.1. Along with lower genera quasitrees, they appear in the definition of the Bollobás–Riordan polynomial of the cellulation [CKS11], which is known to fit in the general framework of Tutte polynomials of matroids [MS18].
Here we make a link to the work of [LLY25].
@@ 12.1. Rooted spanning connected subgraphs
12.1.1. Random $g$-quasitrees
The following example is inspired by the recent preprint [LLY25], which appeared on the arXiv during the first revision of our paper. The first-named author thanks Wai Yeung Lam for a helpful conversation clarifying their work.
As in Section 9.1, let $\mathsf{G}$ denote the $1$-skeleton of a cellular decomposition $X$ of a compact oriented surface $\Sigma$ of genus $g\geqslant 1$. In this setting, a discrete analogue of the period matrix of $\Sigma$ was defined by Mercat (see [Mer07] for a survey). Unlike in the classical case where the period matrix is a $g\times g$ complex matrix, the discrete period matrix is a real $2g\times 2g$ matrix, which we denote by $L$. The reason for this difference is that discrete holomorphic forms are best defined through their real and imaginary parts which are conjugate discrete harmonic forms on the graph and its Poincaré dual.
The main result of [LLY25], Theorem 1.2, expresses all minors of $L$ as weighted (possibly signed) sums over $k$-quasitrees in the surface graph. In the terminology of these authors, a $k$-quasitree is a subgraph $K\in\mathcal{C}_{k}(\mathsf{G})$ such that the induced map between homology groups $H_{1}(K,\mathbb{Z})\to H_{1}(X,\mathbb{Z})$ is injective.
Our observation is that [LLY25, Theorem 1.2] can be interpreted as an equality of bilinear forms on $^{k}\Omega^{1}(\mathsf{G})$: one defined via the period matrix, and the other via a ratio of graph polynomials. Interestingly, the identities of our Proposition 8.2, taken for all $k\in\{1,\ldots,2g\}$, yield the corresponding equalities of quadratic forms, which by polarization encode the same information.
More precisely, this relationship becomes apparent for example by a reformulation of [LLY25, Proposition 3.3]. In order to state it, let us introduce the matrix
$\Omega=\begin{pmatrix}0&I_{g}\\ -I_{g}&0\end{pmatrix}\,.$
Now, let $(c_{1},\ldots,c_{2g})$ be a symplectic basis of the first homology group of $X$, that is a family of simple oriented loops in $C_{1}(\mathsf{G},\mathbb{Z})$ which form an integral basis of $H_{1}(X,\mathbb{Z})$ in which the matrix of the intersection form is $\Omega$. Let $(m_{i}:1\leqslant i\leqslant 2g)$ be the dual basis, that is the collection of $1$-forms such that $m_{i}(c_{j})=\delta_{i,j}$ for all $i,j\in\{1,\ldots,2g\}$. Then [LLY25, Proposition 3.3] may be reformulated as an identity between a Gram matrix and the period matrix $L$:
(114) $\left(\langle m_{i},{\sf\Pi}^{\ker d^{*}}m_{j}\rangle\right)_{1\leqslant i,j\leqslant 2g}=\Omega L\,.$
In our setup, computing principal minors of $\Omega L$ thus corresponds to applying Proposition 8.2 to subspaces of the form $H=\mathop{\mathrm{im}}d\oplus{\mathrm{Vect}}(m_{i_{1}},\ldots,m_{i_{k}})$, for all $k\in\{1,\ldots,2g\}$ and all index sets $\{i_{1}<\cdots<i_{k}\}\subseteq\{1,\ldots,2g\}$. Moreover, by Theorem 8.1, the support of the determinantal probability measures corresponding to these subspaces are bases of a matroid. The special case of [LLY25, Corollary 1.3] corresponds to the case $k=1$ of (3).
Although identities from both papers coincide in this setting, they arise from distinct perspectives. The work of [LLY25] is motivated by Lam’s earlier research on circle patterns and discrete complex structures, as well as the analogy between the period matrix and the response matrix of electrical networks, allowing the authors to resolve a question of Kenyon. Our approach, although also geometrically inspired, arose from the study of determinantal probability measures and applies in greater generality – specializing here to the context of surface graphs.
This crossing of perspectives yields new insight. In particular, Section 7.1 of [LLY25] suggests a geometric refinement of harmonic 1-forms that we had not mentioned in the first version of our paper. Indeed it points to the importance of Lagrangian subspaces (which are of dimension $g$) within the space of harmonic 1-forms (which has dimension $2g$). These are the subspaces of harmonic $1$-forms obtained from the span of $g$ $1$-forms dual to a set of disjoint, homologically independent cycles. This highlights a distinguished family of determinant probability measures on random $g$-quasitrees associated with a surface discretization. These are parametrized by the compact manifold (of dimension $g(g+1)/2$) of all Lagrangian subspaces, which is called the Lagrangian Grassmannian. See Figure 7 for an example.
Figure 7. The $2$-core of a determinantal $2$-quasitree associated with $1$-forms dual to the two meridians.
These measures are especially appealing since $g$-quasitrees are Poincaré dual to $g$-quasitrees on a surface of genus $g$. One may ask whether there exists a canonical choice of Lagrangian subspace yielding a determinantal measure that is self-dual under Poincaré duality. Indeed from the point of view of statistical mechanics, and in particular the Fortuin–Kasteleyn model, such a self-dual model may be critical. If it existed, such a measure could then be viewed as the natural random analogue, on a compact oriented surface of genus $g$, of the uniform spanning tree.
@@ 14.1. Bicircular case
14.1.1. Linear representation of the bicircular matroid
For more details about the content of this section, see [Ken11, Kas15, KL22, KL25].
We consider a connection, that is the data, for each half-edge $(\underline{e},e)$ or $(e,\overline{e})$, of a unitary complex number $h_{\underline{e},e}$ or $h_{\overline{e},e}$. We set $h_{e,\underline{e}}=h_{\underline{e},e}^{-1}$ and $h_{e,\overline{e}}=h_{\overline{e},e}^{-1}$. Given a connection, we define a twisted coboundary map $d_{h}:\Omega^{0}(\mathsf{G})\to\Omega^{1}(\mathsf{G})$ by
$d_{h}f(e)=h_{e,\overline{e}}f(\overline{e})-h_{e,\underline{e}}f(\underline{e}),$
and let $d_{h}^{*}$ be the adjoint linear map. The twisted Laplacian is defined by $\Delta_{h}=d_{h}^{*}d_{h}$. Then for all $f\in\Omega^{0}(\mathsf{G})$, and all $v\in{\sf V}$, we have $\Delta_{h}f(v)=\sum_{e\in{\sf E}:\overline{e}=v}x_{e}(f(v)-h_{v,\underline{e}}f(\underline{e}))$ where $h_{v,\underline{e}}=h_{v,e}h_{e,\underline{e}}$. We still denote by $\Delta_{h}$ the matrix of this operator in the canonical basis of $\Omega^{0}$ indexed by vertices.
By a theorem of Forman [For93], rediscovered and extended to quaternions by Kenyon [Ken11] (see [KL20] for a short proof and references therein), for each $r$, the $r\times r$ principal minors of $\Delta_{h}$ are a weighted sum over bases of the $r$-truncation of the bicircular matroid. We denote by $\mathcal{B}_{-k}(\mathsf{G})$ the set of bases of this matroid, with $r+k=|{\sf V}|$. In particular, $\mathcal{B}_{0}(\mathsf{G})=\mathcal{U}(\mathsf{G})$.
Theorem 14.1 (Forman, Kenyon).
Let $W\subseteq{\sf V}$ be a subset of size $k$. Then
(144) $\det(\Delta_{h})_{\widehat{W}}^{\widehat{W}}=\sum_{F\in\mathcal{B}_{-k}(\mathsf{G})}\prod_{c\,\text{\rm cycle}}|1-\mathrm{hol}_{h}(c)|^{2}\;\underline{x}^{F}\mathbbm{1}_{\{F\,\text{\rm separates points of}\;W\sqcup\{\text{\rm cycles}\}\}}\,$
where for each $F\in\mathcal{B}_{-k}(\mathsf{G})$, the product is over the set of cycles, each with a fixed chosen orientation.
@@ 14.1.2. Determinantal cycle-rooted spanning forests
This twisted matrix-tree formula (Theorem 14.1 in the case $W=\varnothing$) implies the existence of a determinantal probability measure on $\mathcal{U}(\mathsf{G})$ according to Proposition 5.3, a fact first proved by Kenyon [Ken11] who rediscovered Forman’s result, and extended it to the quaternion case. See Figure 10 for a sample. This is the analogue, in this twisted setting, of the Burton–Pemantle theorem (Proposition 6.1) and its first proof.
Theorem 14.2 (Kenyon).
The probability distribution on $\mathcal{U}(\mathsf{G})$ which assigns any $F$ a weight proportional to
(145) $w(F)\,\underline{x}^{F}=\underline{x}^{F}\prod_{c\,\text{\rm cycle}}|1-\mathrm{hol}_{h}(c)|^{2}\,,$
is determinantal associated with the subspace $\mathop{\mathrm{im}}d_{h}$ of $\Omega^{1}(\mathsf{G})$ in the orthonormal basis $(\frac{e^{\star}}{\sqrt{x_{e}}}:e\in{\sf E}^{+})$.
@@
We see from (145) that the support of the measure is a certain subset of $\mathcal{U}(\mathsf{G})$ corresponding to those $F$ all of whose cycles have non-trivial holonomy. This defines a submatroid of the bicircular matroid.
One sees from (145) that the determinantal distribution on $\mathcal{U}(\mathsf{G})$ only depends on the holonomy representation $\rho$ of the fundamental group defined by the connection $h$ (see e.g. [CK23] for a review of these notions).
14.1.3. Symanzik polynomials in the bicircular case
Let us define a family of multivariate polynomials in the Euclidean weights $x_{e}$ associated to the linear representation of the bicircular matroid we just introduced, and more precisely to its holonomy representation $\rho$.
First, we define the weighted basis generating polynomial for the bicircular matroid ‘measured’ by its representing map $d_{h}^{*}$. It reads
(146) $\mathsf{B}^{\rho}(\underline{x})=\sum_{F\in\mathcal{U}(\mathsf{G})}\prod_{c\,\text{cycle}}|1-\mathrm{hol}_{h}(c)|^{2}\underline{x}^{F}\,.$
This is the special case of the polynomial $\mathsf{B}(\underline{x})$ defined in (130) in Section 13. From Theorem 14.1, we have
(147) $\mathsf{B}^{\rho}(\underline{x})=\det\Delta_{h}\,.$
Next, we define a family of polynomials indexed by positive integers $k$, corresponding to weighted basis generating polynomials of truncations of the bicircular matroid. The bases of the $r$-truncation of the bicircular matroid are multi-type spanning forests, that is spanning subgraphs such that each connected component is either a tree or a unicycle (a graph with first Betti number equal to $1$), and such that there are exactly $k=|{\sf V}|-r$ trees (the number of unicycles is not fixed) [^12] Note that such multi-type spanning forests have been considered by Kenyon [Ken11, Ken19] and more recently in a different context motivated by applications by [FB25, JABT24]). Recall that we denote $\mathcal{B}_{-k}(\mathsf{G})$ this set of bases, and further note that $\mathcal{F}_{k-1}(\mathsf{G})\subset\mathcal{B}_{-k}(\mathsf{G})$.
For all positiver integer $k$ and all $k$-tuple ${\sf q}=(q_{1},\ldots,q_{k})$ of elements of $\Omega^{0}(\mathsf{G})$, we set
(148) $\mathsf{A}^{\rho}_{k}({\sf q},\underline{x})=\sum_{F\in\mathcal{B}_{-k}(\mathsf{G})}|\det\left(q_{i}({\sf V}(T_{j}))\right)_{1\leqslant i,j\leqslant k}|^{2}\;\prod_{c\,\text{cycle}}|1-\mathrm{hol}_{h}(c)|^{2}\;\underline{x}^{F}\,,$
where for each $F\in\mathcal{B}_{-k}(\mathsf{G})$, we write $\{T_{1},\ldots,T_{k}\}$ for its connected components which are trees, and the product is over the simple cycles of the unicyclic connected components. These polynomials are the analogues of the Symanzik polynomials ${\mathsf{A}}_{k}({\sf c},\underline{x})$ defined in (70) of Section 8.
14.1.4. Twisted Green function minors
Let us assume that there is at least one cycle $c$ in $\mathsf{G}$ such that $\mathrm{hol}_{h}(c)\neq 1$ (this is equivalent to $\rho\neq 1$). Otherwise, we are back in the circular case of Section 8. This implies that $\Delta_{h}$ is invertible, and we let $G_{h}$ be the inverse of $\Delta_{h}$. The following proposition is the analogue, in the bicircular matroid case, of Proposition 10.5 which holds for the circular matroid case.
Proposition 14.3.
For all ${\sf q}=(q_{1},\ldots,q_{k})$ $k$-tuple of elements of $\Omega^{0}(\mathsf{G})$, we have
(149) $\frac{\mathsf{A}^{\rho}_{k}({\sf q},\underline{x})}{\mathsf{B}^{\rho}(\underline{x})}=\det\left(\langle q_{i},G_{h}q_{j}\rangle\right)_{1\leqslant i,j\leqslant k}\,.$
[…]
Proof.
We first note that each side of the equation we want to prove is a quadratic form in $q_{1}\wedge\ldots\wedge q_{k}$. Hence it is enough to prove it for an orthonormal basis of $^{k}\Omega^{0}(\mathsf{G})$, and then conclude by bilinearity. We will thus prove the formula for each collection of disjoint $k$ vertices $v_{i}$, letting $q_{i}=\mathbbm{1}_{v_{i}}$.
In that case, the right-hand side is the principal minor of $G_{h}$ indexed by $W=\{v_{1},\ldots,v_{k}\}$, which by Jacobi’s complementary minors formula is the quotient of $\det(\Delta_{h})_{\widehat{W}}^{\widehat{W}}$ by $\det\Delta_{h}$.
Now, we turn to the numerator of the left-hand side. By (148), this is the sum over $F$ of a weight which we now compute. We observe that
$\det\left(q_{i}({\sf V}(T_{j}))\right)_{1\leqslant i,j\leqslant k}$
is nonzero if and only if there exists a bijection $\sigma$ of $\{1,\ldots,k\}$ such that $v_{i}$ belongs to $T_{\sigma(i)}$, in which case, the determinant is the signature of this permutation matrix, equal to ±$\pm 1$. Hence in that case, $\mathsf{A}^{\rho}_{k}({\sf q},\underline{x})=\det(\Delta_{h})_{\widehat{W}}^{\widehat{W}}$. The result now follows by invoking (147). ∎
We note that since $d_{h}^{*}$ is onto, we may write for each $i\in\{1,\ldots,k\}$, that $q_{i}=d_{h}^{*}\phi_{i}$, for $\phi_{i}\in\Omega^{1}(\mathsf{G})$. Then, letting $\varphi=\phi_{1}\wedge\ldots\wedge\phi_{k}$, Equation (149) reads
(150) $^{k}{\sf\Pi}^{\mathop{\mathrm{im}}d_{h}}(\varphi)\rVert^{2}=\frac{\mathsf{A}^{\rho}_{k}({\sf q},\underline{x})}{\mathsf{B}_{\rho}(\underline{x})}\,.$
14.1.5. Determinantal multi-type spanning forests
Given $k$ functions $q_{1},\ldots,q_{k}\in\Omega^{0}(\mathsf{G})$, let
(151) $H_{{\sf q}}=\{d_{h}f:f\in\Omega^{0}(\mathsf{G}),\forall 1\leqslant i\leqslant k,\langle f,q_{i}\rangle=0\}\,.$
For any choice of $\phi_{1},\ldots,\phi_{k}\in\Omega^{1}(\mathsf{G})$ such that $q_{i}=d_{h}^{*}\phi_{i}$ for all $i\in\{1,\ldots,k\}$, we may rewrite this space as
(152) $H_{{\sf q}}=\mathop{\mathrm{im}}d_{h}\cap{\mathrm{Vect}}(\phi_{1},\ldots,\phi_{k})^{\perp}\,.$
The following theorem is the analogue, in the bicircular matroid case, of Theorem 11.3 which holds for the circular matroid case. See Figure 11 for an example.
Figure 11. A determinantal random element of $\mathcal{B}_{-2}(\mathsf{G})$.
Theorem 14.4.
The probability distribution on $\mathcal{B}_{-k}(\mathsf{G})$ associated to the polynomial $\mathsf{A}^{\rho}_{k}({\sf q},\underline{x})$, seen as the generating function of a positive measure, is determinantal associated to the orthogonal projection on $H_{{\sf q}}$ in the canonical orthonormal basis of $\Omega^{1}(\mathsf{G})$ indexed by ${\sf E}^{+}$.
Proof.
We follow the same strategy as in the first proof of Theorem 8.3. Hence we want to apply the second part of Proposition 5.3 and for that we need an expression for the partition function of our probability measure of the form $\underline{x}^{{\sf E}}\det bb^{*}$ where $\ker b=H_{{\mathsf{q}}}$.
For that, we first need to prove an analogue of the second equality in Proposition 8.4. We will do so using Lemma 13.1 in place of (51).
Let us start by defining the map $b$. We follow word for word the construction in the circular matroid case. The only difference is that we do not use $Z_{1}(\mathsf{G},\mathbb{K})$. Instead we replace that space with $Z=({\sf J}^{1}_{\underline{1}})^{-1}\ker d_{h}^{\dagger}\subset\Omega_{1}(\mathsf{G})$. We let $c_{i}$ be the preimages by ${\sf J}^{1}_{\underline{x}}$ of $\phi_{i}$.
The key point is that we will observe how changing the Euclidean weights from $1$ to $\underline{x}$ changes the volume of an arbitrary basis $\mathscr{Z}$ of $Z$ (we have no integral structure here unlike in the circular matroid case). This is where Lemma 13.1 comes into play. We find that
$\lVert\det\mathscr{Z}\lVert^{2}=\sum_{T\in\mathcal{U}(\mathsf{G})}\det(\mathscr{Z}/\mathscr{Z}_{T})^{2}\,\frac{\underline{x}^{T}}{\underline{x}^{\sf E}}\,.$
Let us call $\mathcal{V}(\underline{x)}$ this function of $\underline{x}$. By the same argument as in the proof of Proposition 8.4, we find that $\det(1^{Q}bb^{*}1_{Q})$ is equal to $\mathcal{V}(\underline{x})/\mathcal{V}(\underline{1})$.
But now, combining Proposition 13.6 and (145), we find that
$T\mapsto\frac{w(T)}{(\det\mathscr{Z}/\mathscr{Z}_{T})^{2}}$
is a constant function of $T\in\mathcal{U}(\mathsf{G})$. Hence, the above ratio of volume $V(\underline{x})/V(\underline{1})$ is proportional to $\mathsf{B}_{\rho}(\underline{x})/\underline{x}^{{\sf E}}$ and evaluating it at $\underline{x}=\underline{1}$ where it is equal to $1$, we find that the volume-ratio is thus equal to
$\frac{\mathcal{V}(\underline{x})}{\mathcal{V}(\underline{1})}=\frac{\mathsf{B}_{\rho}(\underline{x})}{\mathsf{B}_{\rho}(\underline{1})\underline{x}^{{\sf E}}}\,.$
Hence, combining (150), and the same Schur lemma (Lemma 5.7) argument as in the proof of Proposition 8.4 we are following step by step, we obtain
(153) $\mathsf{A}^{\rho}_{k}({\mathsf{q}},\underline{x})=\mathsf{B}_{\rho}(\underline{x})\big{\|}\mathchoice{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{1.0pt}{$\bigwedge$}}{\raisebox{0.5pt}{$\scriptstyle\bigwedge$}}{\raisebox{0.2pt}{$\scriptscriptstyle\bigwedge$}}^{k}{\sf\Pi}^{\mathop{\mathrm{im}}d_{h}}(\varphi)\big{\|}^{2}=\mathsf{B}_{\rho}(\underline{1})\underline{x}^{\sf E}\det(bb^{*})\,.$
The result now follows by an application of the second determinantality criterion (second assertion of Proposition 5.3). ∎
14.1.6. An example on the ‘rank $1$ extension’ of the bicircular matroid
@@
Let us consider a unitary rank $1$ connection $h$ on $\mathsf{G}$. As we have just seen, the subspace $\mathop{\mathrm{im}}d_{h}$ of $\Omega^{1}(\mathsf{G})$ defines a determinantal probability measure on $\mathcal{U}(\mathsf{G})$. And by considering subspaces $H\subseteq\mathop{\mathrm{im}}d_{h}$, we defined determinantal multi-type spanning forests. Let us now take a look at what happens if we take a subspace $H$ containing $\mathop{\mathrm{im}}d_{h}$, in the simplest case where the codimension of $\mathop{\mathrm{im}}d_{h}$ in $H$ is $1$.
Hence, let us consider a $1$-form $\theta$ outside $\mathop{\mathrm{im}}d_{h}$ and consider the subspace
$H_{\theta}=\mathop{\mathrm{im}}d_{h}\oplus\mathbb{C}\theta\subseteq\Omega^{1}(\mathsf{G})\,.$
@@
By Theorem 13.13, the corresponding determinantal measure is supported on the set of spanning subgraphs $K$ with connected components $K_{1},\ldots,K_{m}$, all of which have first Betti number equal to $1$, but $K_{1}$, which has first Betti number equal to $2$.